# Gravitational Waves on Kerr Black Holes I: Reconstruction of Linearized Metric Perturbations

### v2

@@ F.1 Mapping the angular ODE to the confluent Heun equation
This equation has two regular singular points at $u=\pm 1$ and an irregular singular point at $u=\infty$. This means that via a change of variable and a field redefinition, it can be mapped to the canonical form (E.1) of the confluent Heun equation, for some parameters $\gamma$, $\delta$, $\epsilon$, $\alpha$, and $q$. To put it into this canonical form, we first perform the redefinition [^25] The $+1$ in the exponent has no effect on what follows and simply scales the function by a constant. We add it here to match Eq. (2.6) and prevent factors of $e^{a\omega}$ from appearing in many equations.
@@ F.2 Boundary conditions and quantization of the mode spectrum
The variables $z_{+}$ and $z_{-}$ are related by $u\to-u$, or $\theta\to\pi-\theta$. In other words, they are related by a reflection through the equatorial plane $\theta=\frac{\pi}{2}$. Since the Kerr metric (1.1) enjoys an equatorial reflection symmetry, it follows that the choice of either $z_{+}$ or $z_{-}$ is arbitrary, so from now on, we consider only the mapping to $z_{+}$ and drop the $+$ subscript.
@@
In summary, the boundary conditions (and our choices) fix the angular modes $S_{\omega\ell m}^{(s)}(z)$ to take the form (F.2) with $\mu_{1}$ and $\mu_{2}$ given in Eq. (F.12), $\mu_{3}=a\omega$, and $H(z)=\operatorname{HeunC}(q,\alpha,\gamma,\delta,\epsilon;z)$ a confluent Heun function with parameters given with a $+$ in Eq. (F.9). Moreover, the accessory parameter $q$, and hence the separation constant $\lambda_{\omega\ell m}^{(s)}$, are quantized via the condition (F.13).
@@ G.1 Mapping the radial ODE to the confluent Heun equation
+$+\mathopen{}\mathclose{{}\left\{\xi_{3}^{2}+\omega^{2}+\frac{1}{\mathopen{}\mathclose{{}\left(r-r_{+}}\right)^{2}}\mathopen{}\mathclose{{}\left[\xi_{1}-ic_{+}\mathopen{}\mathclose{{}\left(\omega-m\Omega_{+}}\right)}\right]\mathopen{}\mathclose{{}\left[\xi_{1}+s+ic_{+}\mathopen{}\mathclose{{}\left(\omega-m\Omega_{+}}\right)}\right]}\right.$

### v3

@@ F.1 Mapping the angular ODE to the confluent Heun equation
This equation has two regular singular points at $u=\pm 1$ and an irregular singular point at $u=\infty$. This means that via a change of variable and a field redefinition, it can be mapped to the canonical form (E.1) of the confluent Heun equation, for some parameters $\gamma$, $\delta$, $\epsilon$, $\alpha$, and $q$. To put it into this canonical form, we first perform the redefinition [^25] The +1+1 in the exponent has no effect on what follows and simply scales the function by a constant. We add it here to match Eq. (2.6) and prevent factors of $e^{a\omega}$ from appearing in many equations.
@@ F.2 Boundary conditions and quantization of the mode spectrum
The variables $z_{+}$ and $z_{-}$ are related by $u\to-u$, or $\theta\to\pi-\theta$. In other words, they are related by a reflection through the equatorial plane $\theta=\frac{\pi}{2}$. Since the Kerr metric (1.1) enjoys an equatorial reflection symmetry, it follows that the choice of either $z_{+}$ or $z_{-}$ is arbitrary, so from now on, we consider only the mapping to $z_{+}$ and drop the ++ subscript.
@@
In summary, the boundary conditions (and our choices) fix the angular modes $S_{\omega\ell m}^{(s)}(z)$ to take the form (F.2) with $\mu_{1}$ and $\mu_{2}$ given in Eq. (F.12), $\mu_{3}=a\omega$, and $H(z)=\operatorname{HeunC}(q,\alpha,\gamma,\delta,\epsilon;z)$ a confluent Heun function with parameters given with a ++ in Eq. (F.9). Moreover, the accessory parameter $q$, and hence the separation constant $\lambda_{\omega\ell m}^{(s)}$, are quantized via the condition (F.13).
@@ G.1 Mapping the radial ODE to the confluent Heun equation
$+\mathopen{}\mathclose{{\left\{\xi_{3}^{2}+\omega^{2}+\frac{1}{\mathopen{}\mathclose{{\left(r-r_{+}}}\right)^{2}}\mathopen{}\mathclose{{\left[\xi_{1}-ic_{+}\mathopen{}\mathclose{{\left(\omega-m\Omega_{+}}}\right)}}\right]\mathopen{}\mathclose{{\left[\xi_{1}+s+ic_{+}\mathopen{}\mathclose{{\left(\omega-m\Omega_{+}}}\right)}}\right]}}\right$.